# Forecasting When to Forecast: Accelerating Diffusion Models with Confidence-Gated Taylor

### v1

@@ 3 Method
Figure 2: The framework of our method. At each timestep $t$, we first compute the actual output of the first block and simultaneously predict it using Taylor expansion. The prediction error is then evaluated. If the error is below a threshold $\epsilon$, indicating that the Taylor prediction is reliable, we use it to approximate the last block feature-skipping the computation of the remaining $B{-}1$ blocks to accelerate inference. Otherwise, we fall back to full computation of all blocks. This process is repeated for each subsequent timestep.
@@ 3.1 Preliminary
where $\mathcal{F}(x_{t}^{b})$ denotes features at fully activated timesteps and $\Delta^{i}\mathcal{F}(x_{t}^{b})$ represents the $i$-th order finite difference. To avoid explicit computation, they approximated higher-order derivatives using finite differences. Then the features for the $t-k$ steps are predicted as:
¯$\overline{\mathcal{F}}{(x_{t-k}^{b})}=\mathcal{F}(x_{t}^{b})+\sum_{i=1}^{O}\frac{\Delta^{i}\mathcal{F}(x_{t}^{b})}{i!\cdot N^{i}}(-k)^{i}.$ (5)
@@ 3.2 Last Block Forecast
Then, we estimate future outputs for the $t-k$ step using a Taylor expansion:
@@ 3.3 Prediction Confidence Gating
We discover that Transformer-based architectures like DiT, the blocks exhibit strong sequential dependencies—meaning the quality of early block predictions often reflects the predictability of later blocks. Motivated by this insight, we introduce a lightweight proxy estimation method, Prediction Confidence Gating (PCG), which uses the prediction error of the first block to estimate the reliability of Taylor-based prediction at the current timestep (Figure 2). Specifically, at each timestep $t-k$, we apply Taylor expansion to predict the first Transformer block output, using the same expansion form defined in Eq. 7, but applied at the first block:
@@ 4.1 Experiment Settings
FORA ($N=3$[28] 8.58 0.197 2.17x 3.55 6.36 229.02
ToCa ($N=3$[50] 10.23 0.216 1.98x 2.87 4.76 235.21
DuCa ($N=3$[49] 9.58 0.208 2.06x 2.88 4.66 233.37
Taylorseer ($N=3,O=3$[15] 8.56 0.248 1.73x 2.34 4.69 238.42

### v2

@@ 3 Method
Figure 2: The framework of our method. At each timestep $t$, we first compute the actual output of the first block and simultaneously predict it using Taylor expansion. The prediction error is then evaluated. If the error is below a threshold $\epsilon$, indicating that the Taylor prediction is reliable, we use it to approximate the last block feature-skipping the computation of the remaining B1B{-}1 blocks to accelerate inference. Otherwise, we fall back to full computation of all blocks. This process is repeated for each subsequent timestep.
@@ 3.1 Preliminary
where $\mathcal{F}(x_{t}^{b})$ denotes features at fully activated timesteps and $\Delta^{i}\mathcal{F}(x_{t}^{b})$ represents the $i$-th order finite difference. To avoid explicit computation, they approximated higher-order derivatives using finite differences. Then the features for the tkt-k steps are predicted as:
$\overline{\mathcal{F}}{(x_{t-k}^{b})}=\mathcal{F}(x_{t}^{b})+\sum_{i=1}^{O}\frac{\Delta^{i}\mathcal{F}(x_{t}^{b})}{i!\cdot N^{i}}(-k)^{i}.$ (5)
@@ 3.2 Last Block Forecast
Then, we estimate future outputs for the tkt-k step using a Taylor expansion:
@@ 3.3 Prediction Confidence Gating
We discover that Transformer-based architectures like DiT, the blocks exhibit strong sequential dependencies—meaning the quality of early block predictions often reflects the predictability of later blocks. Motivated by this insight, we introduce a lightweight proxy estimation method, Prediction Confidence Gating (PCG), which uses the prediction error of the first block to estimate the reliability of Taylor-based prediction at the current timestep (Figure 2). Specifically, at each timestep tkt-k, we apply Taylor expansion to predict the first Transformer block output, using the same expansion form defined in Eq. 7, but applied at the first block:
@@ 4.1 Experiment Settings
FORA (N=3N=3[28] 8.58 0.197 2.17x 3.55 6.36 229.02
ToCa (N=3N=3[50] 10.23 0.216 1.98x 2.87 4.76 235.21
DuCa (N=3N=3[49] 9.58 0.208 2.06x 2.88 4.66 233.37
Taylorseer (N=3,O=3N=3,O=3[15] 8.56 0.248 1.73x 2.34 4.69 238.42